OpenStudy (anonymous):

find an equation of the tangent line to the graph of the function at the given point. f(x)e^3 1nx,(1,0)

OpenStudy (anonymous):

differentiate and sub in x =1 and y = 0

OpenStudy (campbell_st):

can you find the 1st derivative

OpenStudy (anonymous):

f(x) = e^3x ^ ln(x) .... product rule is needed dy/dx = dv/dx * u + du/dx * v let u = e^3x and v = ln(x) ==> therefore du/dx = 3e^3x and dv/dx = 1/x so therefore dy/dx = (e^3x)/x + 3e^3x*lnx so e^3 + 3e^3*ln1 should equal the gradient

OpenStudy (anonymous):

@Iamgmg90 your amazing thanks! (:

OpenStudy (anonymous):

of the tangent at that point, which is equal to m use the formula y_2 - y_1 = m(x_2 - x_1) for x = 1 and y = 0 to calculate the formula, you must be given another coordinate, or be able to infer one

OpenStudy (campbell_st):

just a quick question is the function 1. \[f(x) = e^{3\ln(x)}\] or 2. \[f(x) = e^{3x^{\ln(x)}}\] as there will be a considerable difference in the derivative.

OpenStudy (anonymous):

@campbell_st http://learn.flvs.net/educator/temp/atucker69/2579/ChapterFive/5.04.pdf put that in the search bar its number 56

OpenStudy (campbell_st):

its just that \[f(x) = e^{3\ln(x)} = e^{\ln(x^3)} = x^3\] using the log laws.... just a thought... so the question is \[f(x) = e^3 \ln(x) \] so e^3 is a constant so the derivative is \[f'(x) = \frac{e^3}{x}\] as I said... just a thought...

OpenStudy (anonymous):

@campbell_st thank you (: